OpenStudy (xjason):

The length of each side of a square increases by 2.5 inches to form a new square with a perimeter of 70 inches. The length of each side of the original square was _________ inches.

OpenStudy (data_lg2):

Perimeter is the total length of the sides of a polygon. A square has 4 equal sides, so the formula for its perimeter is \(\sf P=L1+L2 +L3 + L4= 4L\)  From your question we have square which was increased by 2.5 and will have a perimeter of 70 inches, so your equation will look like this: \(\sf 70 = 4 \times 2.5 \times L\) Now you have to find the value of L by simplifying the equation: \(\sf 70 = 4 \times 2.5 \times L\)

OpenStudy (data_lg2):

can you try solving it? If you get stuck, don't hesitate to ask (:

OpenStudy (xjason):

i dont get this

OpenStudy (data_lg2):

which part you did not get?

OpenStudy (xjason):

l

OpenStudy (xjason):

L

OpenStudy (data_lg2):

I use the variable "L" as representation for the side of the square...

OpenStudy (xjason):

oh ok

OpenStudy (data_lg2):

\(\sf P=side\ length + side\ length + side\ length + side\ length = 4 \times side\ length \) we have to multiply 2.5 because the square is increased, and the perimeter becomes 70: \(\sf 70= 4 \times side\ length \times 2.5\)

OpenStudy (data_lg2):

Maybe we can do solve it together... let's do the right side first. What is \(\sf 2.5 \times 4\)?

OpenStudy (xjason):

10

OpenStudy (data_lg2):

right.. so now you have: \(\sf 70= 10 \times side\ length\) Now you have to divide both sides by 10, what will you get?

OpenStudy (xjason):

wait so divide 70 and side lengths by 10

OpenStudy (data_lg2):

yes.... divide 70 by 10 and divide 10 side lengths by 10

OpenStudy (xjason):

ok

OpenStudy (xjason):

let me do it

OpenStudy (data_lg2):

okay, tell me what you will get ^_^

OpenStudy (xjason):

im not sure how you divide 10 side lengths by 10

OpenStudy (data_lg2):

the 10s will get cancelled out ^_^

OpenStudy (data_lg2):